# RLinf: Flexible and Efficient Large-scale Reinforcement Learning via Macro-to-Micro Flow Transformation

### v1

@@ 5.1 Experimental Setup
    Minibatch Early-Stop: We discarde minibatches with too large importance ratio to stabilize training.
We use a rule-based reward funtion, which will return 5 if the final boxed/numeric answer is correct else -5.
For embodied models, we use both PPO and GRPO for ManiSkill+OpenVLA and Libero+OpenVLA-OFT.
5.2 End-to-End Performance
Reasoning RL Training.   Our evaluation demonstrates that RLinf consistently outperforms baseline systems across all model scales. As shown in Figures 8(a), 8(b), and 8(c), which present RLHF throughput results for GRPO on Qwen2.5 models at 1.5B, 7B, and 32B parameters, RLinf achieves a speedup ranging from 1.10× to 1.58× compared to veRL.
Moreover, Figure 8 reveals that veRL exhibits suboptimal performance scaling in multi-GPU environments. This limitation is primarily attributed to two factors: (1) its log probability computation becomes a computational bottleneck, as will be discussed in detail in Section  5.3; (2) the unoptimized rollout engine results in excessive peak GPU memory usage, necessitating a reduction in the memory allocated for the rollout engine’s KV cache. This constraint further diminishes its scalability.
As previously noted in Section 2.2, the dynamic and long-tailed nature of response generation results in GPU underutilization, a problem that becomes more pronounced in long-context RL training. Figure 10 compares performance between collocated and disaggregated modes under a context length of 28,672 and group size 8. The disaggregated mode yields a speedup of 1.17× to 1.21× over the collocated mode.
Figure 10: Throughput of RLinf under different placement modes and model size settings.
Embodied RL Training.   Figure 9 presents the end-to-end throughput of RLinf and baseline methods in embodied AI training under different cluster sizes and placement strategies, with ManiSkill running on GPU and LIBERO on CPU.
For ManiSkill, RL4VLA is integrated under a disaggregated mode configuration. As shown in 9(a), RLinf achieves a 1.88× speedup in hybrid mode on 8 GPUs. When scaling to 16 and 32 GPUs, overhead from model loading/offloading and state switching increases; however, the hybrid mode mitigates this cost through pipelined execution that overlaps computation and communication, still delivering a 1.61× to 1.69× speedup.
Although the hybrid mode yields significant improvements in ManiSkill, 9(b) shows that the collocated mode performs better in LIBERO, achieving a 1.25× to 2.13× speedup. Since LIBERO is a CPU-bound task, the training process is bottlenecked by CPU-based rollout. The collocated mode allocates all resources to the rollout stage, resulting in higher performance in this scenario.
5.3 Performance Breakdown
Figure 11 illustrates the latency breakdown for training the Qwen2.5-7B model, corresponding to the setup in 8(b). The performance of veRL is constrained by two main factors: first, its unoptimized rollout engine requires a reduction in KV cache memory allocation, which prolongs the rollout process; second, veRL exhibits longer inference times, further reducing its overall throughput.
Figure 12 presents a latency breakdown and comparison between collocated and disaggregated modes for the 7B model, complementing the results in Figure 10. It can be observed that despite allocating only 40 out of 64 GPUs to the rollout stage in disaggregated mode, the end-to-end rollout time increases by merely 14%. Moreover, both the inference and training stages—once their rollout results are ready—can execute concurrently with the rollout of other requests. This parallelization leads to a reduction in end-to-end iteration time and higher overall throughput.
Figure 13 presents a detailed performance breakdown of LIBERO, complementing the results shown in 9(b). RLinf demonstrates superior training efficiency in both hybrid and collocated modes. The improved rollout performance can be attributed to two key optimizations: (1) the elimination of redundant environment initialization during the rollout stage, and (2) the use of a single forward pass to compute both action and log probability, achieved with a modest increase in memory usage.
Figure 11: Latency breakdown of Qwen2.5 7B model training in 8(b).
Figure 12: Latency breakdown and comparison for collocated and disaggregated modes on the 7B model in Figure 10. The width of bar is proportional to the number of GPUs. Note that under the disaggregated mode, the reward computation time is included within the rollout phase and is therefore not shown separately.
Figure 13: Latency breakdown of LIBERO in 9(b).
5.4 Algorithm Performance
We conducted comprehensive evaluation across various benchmarks for different modalities. Our benchmarks primarily consist of the following two categories:
@@ 5.4 Algorithm Performance
Table 4: Evaluation scores of RLinf-math-1.5B/7B and open-source models.
Model AIME 24 AIME 25 GPQA-diamond Average
@@
DeepSeek-R1-Distill-Qwen-1.5B (base model) 28.33 24.90 27.45 26.89
DeepMath-1.5B 37.80 30.42 32.11 33.44
DeepScaleR-1.5B-Preview 40.41 30.93 27.54 32.96
AReaL-1.5B-Preview-Stage-3 40.73 31.56 28.10 33.46
AReaL-1.5B-retrain* 44.42 34.27 33.81 37.50
FastCuRL-1.5B-V3 43.65 32.49 35.00 37.05
RLinf-math-1.5B 48.44 35.63 38.46 40.84
@@
DeepSeek-R1-Distill-Qwen-7B (base model) 54.90 40.20 45.48 46.86
AReaL-boba-RL-7B 61.66 49.38 46.93 52.66
Skywork-OR1-7B 66.87 52.49 44.43 54.60
Polaris-7B-Preview 68.55 51.24 43.88 54.56
AceMath-RL-Nemotron-7B 67.30 55.00 45.57 55.96
RLinf-math-7B 68.33 52.19 48.18 56.23
Table 5: Base Model Mapping For Embodiment RL Models
Model Name Base Model
RLinf-OpenVLA-GRPO-ManiSkill3-25ood
gen-robot/openvla-7b-rlvla-warmup
RLinf-OpenVLA-PPO-ManiSkill3-25ood
RLinf-OpenVLAOFT-GRPO-ManiSkill3-25ood
RLinf/Openvla-oft-SFT-libero10-trajall
and RLinf/RLinf-OpenVLAOFT-ManiSkill-Base-Lora
RLinf-OpenVLAOFT-PPO-ManiSkill3-25ood
RLinf-OpenVLAOFT-GRPO-LIBERO-spatial Haozhan72/Openvla-oft-SFT-libero-spatial-traj1
RLinf-OpenVLAOFT-GRPO-LIBERO-object Haozhan72/Openvla-oft-SFT-libero-object-traj1
RLinf-OpenVLAOFT-GRPO-LIBERO-goal Haozhan72/Openvla-oft-SFT-libero-goal-traj1
RLinf-OpenVLAOFT-GRPO-LIBERO-long Haozhan72/Openvla-oft-SFT-libero-long-traj1
Table 6: OpenVLA and OpenVLA-OFT models success rate results on ManiSkill3
Model Algorithm Vision Semantic Position Average
gen-robot/openvla-7b-rlvla-rl PPO 76.6% 75.4% 77.6% 76.1%
RLinf-OpenVLAOFT-GRPO-ManiSkill3-25ood GRPO 84.6% 51.6% 42.9% 61.5%
RLinf-OpenVLAOFT-PPO-ManiSkill3-25ood PPO 80.5% 56.6% 56.1% 64.5%
RLinf-OpenVLA-GRPO-ManiSkill3-25ood GRPO 74.7% 74.4% 81.6% 75.5%
RLinf-OpenVLA-PPO-ManiSkill3-25ood PPO 82.0% 80.6% 89.3% 82.2%
  • $\bullet$
    Math Benchmark: AIME 24, AIME 25, GPQA-diamond.
  • $\bullet$
    Embody Benchmark: LIBERO, as well as custom pick-and-place tasks on ManiSkill with out-of-distribution (OOD) vision, semantic, and position challenges.
Reasoning RL Training.   Using RLinf, we applied reinforcement learning to both DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B, presenting two new models: RLinf-math-1.5B and RLinf-math-7B. These models achieve best performance across multiple benchmarks among models of comparable scale, as shown in Table 4. RLinf-math-1.5B outperforms all baseline models across all benchmarks, improving by up to 20 points over its base model, DeepSeek-R1-Distill-Qwen-1.5B. RLinf-math-7B achieves the highest performance on GPQA-diamond and leads in average benchmark scores among all compared models.
Table 7: OpenVLA-OFT model results on LIBERO
Model LIBERO
Spatial Object Goal Long Avg.
Octo 78.9 85.7 84.6 51.1 75.1
OpenVLA 84.7 88.4 79.2 53.7 76.5
Nora 92.2 95.4 89.4 74.6 87.9
$\pi_{0}$+FAST 96.4 96.8 88.6 60.2 85.5
$\pi_{0}$ 96.8 98.8 95.8 85.2 94.2
UniVLA 96.5 96.8 95.6 92.0 95.2
OpenVLA-OFT One-Trajectory SFT
Baseline 56.5 25.6 45.6 9.7 34.4
RLinf 99.0 99.0 99.0 94.4 97.9
$\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\Delta$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+42.5$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+73.4$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+53.4$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+84.7$} $\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}+63.5$}
Embodied RL Training.   To achieve reasonable reinforcement learning results in an embodied setting, the base models must exhibit a non-zero success rate. Therefore, we selected different base models for different environments and model types, each of which was initialized from a supervised fine-tuned OpenVLA or OpenVLA-OFT model. The detailed mapping between the final RL models and their base models is provided in Table 5. Note that the RL models are named according to the convention: RLinf-[model type]-[RL algorithm]-[environment].
Table 6 and Table 7 present the evaluation results of our trained models on the ManiSkill and LIBERO benchmarks, respectively. In the ManiSkill environment, the trained models demonstrate strong generalizability under out-of-distribution settings and achieve higher success rates than the baseline RL4VLA [21] checkpoint. On the LIBERO benchmark, our RL models exhibit significant performance improvement, attaining state-of-the-art success rates on all spatial, object, goal, and long tasks.
@@ 6 Related Works
RL Training Frameworks.   RL training frameworks adopt varying system designs to support large-scale alignment tasks [29, 5], typically falling into either task-colocated or task-separated execution modes. For task-colocated systems, DeepSpeed Chat [44] and VERL[39] put all training phases on shared GPUs for simplified orchestration. In contrast, task-separated systems like NeMo-Aligner [38] and OpenRLHF [12] deivde components across devices to improve modularity and scalability. AReal [10] further introduces asynchronous model updates algorithm upon task-separated systems to increase training throughput. RLinf differs from these frameworks by supporting dynamic transitions between colocated and separated execution modes based on runtime workload patterns.
Distributed Training Systems for LLM.   A variety of distributed training frameworks have been developed to support large language models (LLMs) [47, 48, 19], alongside numerous system-level optimizations targeting memory usage [4, 23, 13] and communication efficiency [3, 22, 15]. Megatron-LM [40] adopts hybrid parallelism by combining tensor, pipeline, and data parallelism, significantly enhancing scalability for multi-billion parameter models. DeepSpeed [34] introduces the ZeRO series of optimizations [32, 35, 33], which partition optimizer states, gradients, and activations across devices, drastically reducing per-GPU memory requirements. Distributed reinforcement learning systems share similar challenges with LLM training frameworks in scaling large models across multiple devices under resource constraints. However, RL workflows are inherently more dynamic due to interactive data generation and asynchronous updates, making system-level optimization more complex and challenging.
Dataflow System.   Traditional dataflow systems [46, 14, 25, 6] have played a foundational role in scaling large-scale data processing and machine learning workloads. These systems rely on static task graphs and centralized scheduling, making them highly efficient for batch or streaming jobs with predictable structures. In contrast, reinforcement learning training pipelines typically involve dynamic task graphs with asynchronous components such as data collection, policy updates, and training. These workloads demand greater flexibility in scheduling and coordination than traditional systems provide. Modern frameworks like Ray [24] address this gap with actor-based execution and decentralized control, making them well-suited for orchestrating complex RL workflows.

### v2

@@ 5 Evaluation
    The models trained with RLinf achieve SOTA or better benchmark scores after RL fine-tuning.
Figure 8: GRPO training throughput of Qwen2.5 on RLinf and veRL under different cluster scales and model sizes.
5.1 End-to-End Experiments
The end-to-end experiments run on a 32-node cluster, where each node is equipped with 8 NVIDIA H100-80GB GPUs, 2 Intel Xeon Platinum 8558 CPUs (2.1 GHz, 48 cores), and 2 TB of memory. Intra-node communication uses NVLink, while inter-node communication uses 8 Mellanox ConnectX-7 RDMA NICs per node, each providing 400 Gbps bandwidth with RoCEv2.
5.1.1 Reasoning RL Training
Experimental setup.   We evaluate using Qwen2.5 Models [qwen2025qwen25technicalreport] distilled by DeepSeek-R1,covering sizes from 1.5B to 32B parameters and Qwen3-30B-A3B [yang2025qwen3] which is an MoE model, and using AReaL-boba-Data dataset [arealboba], which integrates multiple standard datasets including DeepScaleR, Open-Reasoner-Zero, Light-R1, etc. We compare RLinf with two state-of-the-art open-source RL systems, i.e., veRL v0.5 [verl] and Slime v0.1 [slime]. To ensure fairness, all systems use SGLang [sglang] for rollouts and Megatron-LM [megatron] for training, with the same parallelism setting (e.g., tensor parallelism configuration). Training speed is reported in tokens/sec, defined as the total number of prompt and response tokens in a global batch divided by the iteration time. All performance results are averaged over 10 training iterations after warm-up.
[…]
Figure 8 also shows that veRL scales poorly as GPU number increases, largely because inference becomes a growing bottleneck, with its proportion of total execution time rising from 15.2% to 19.9% when scaling from 64 to 256 GPUs. In addition, veRL’s unoptimized rollout engine leads to excessive peak memory usage, forcing smaller KV-cache allocations and causing rollout time to decrease only sublinearly.
We also tested spatial mode, but it underperforms veRL by 44.3-68.6% for the 7B mode, because rollout, inference, and training run on disjoint smaller sets of GPUs, slowing both rollout and training. With the long sequence length (i.e., 28672), the training stage also waits longer for the first batch of rollouts to be generated. In addition, the rollout batch size is relatively smaller, reducing the overlapping time.
Figure 9: Latency breakdown of Qwen2.5 7B model training.
Figure 10: PPO training throughput of Qwen2.5 on RLinf and veRL under different cluster scales and model sizes.
Figure 11: Latency breakdown of Qwen2.5 7B with PPO on 32 GPUs. The width of bar presents the number of GPUs.
Figure 12: RL training throughput of Qwen3-30B-A3B on RLinf and Slime under different cluster scales.
Figure 13: Latency breakdown of Qwen3-30B-A3B with GRPO on 32 GPUs. The width of bar is proportional to the number of GPUs.
Qwen2.5 with PPO.   We train PPO with Qwen2.5 as both the actor and critic, using model sizes of 1.5B, 7B, and 14B, scaling from 16 to 256 GPUs. We use a rule-based reward, a maximum sequence length of 12288, and a rollout batch size ranging from 256 to 1024 depending on GPU count. RLinf-Spatial allocates GPUs in a 4:1:1:1:1 ratio for rollout, actor inference, actor training, critic inference, and critic training. Figure 10 presents the results. For the 1.5B model, RLinf-Spatial outperforms veRL by 69.6%, 35.0%, and 35.6% on 16, 32, and 64 GPUs, respectively. In contrast to the GRPO experiment, RLinf-Spatial surpasses RLinf-Temporal by 39.8%, 19.4%, and 18.6%.
For the 7B model, RLinf-Spatial outperforms both veRL and RLinf-Temporal by 38.7–60.7% and 19.0–44.8%, respectively. As shown in Figure 13, RLinf-Temporal exceeds veRL mainly due to speedups in “Others”, which includes context-switch overhead (offload/onload), parameter resharding, advantage computation, and inter-stage synchronization. RLinf-Spatial performs faster than RLinf-Temporal because rollout, inference, and training overlap effectively. Although rollout in spatial mode is 39.3% slower than in temporal mode, inference and training finish quickly once rollout completes. The 14B model shows similar trends with smaller gains, i.e., RLinf-Spatial is 27.2–56.5% and 17.7–25.7% faster than veRL and RLinf-Temporal, respectively.
Qwen3-30B-A3B with GRPO.   We evaluate the MoE model Qwen3-30B-A3B on 32, 64, and 128 GPUs with a rollout batch size of 1536 and sequence length 20480. We disable logprob recomputation, a common configuration in RL, which removes the inference stage. Since Slime is optimized for MoE RL training, we compare against two variants: Slime (spatial mode without pipelining, as Slime does not support pipelining) and Slime-Colocate (temporal mode). As shown in Figure 13, Slime is the slowest because rollout and training run on disjoint GPU sets with no execution overlap. For 32 and 64 GPUs, RLinf-Temporal performs similarly to Slime-Colocate, while RLinf-Spatial (1:1 rollout-to-training GPUs) is 31.2% and 7.2% faster than Slime-Colocate, respectively.
Figure 13 presents the performance breakdown. Spatial mode achieves shorter rollout time because temporal mode suffers from memory contention between rollout and training, forcing much smaller parallelism (i.e., max_running_requests of 128 vs. 256 in SGLang). In addition, RLinf-Spatial achieves good rollout–training overlap. On 128 GPUs, RLinf-Temporal outperforms Slime-Colocate by 3.7% due to better scaling from decoupled system design; both stages run slightly faster. RLinf-Spatial underperforms because rollout and training do not overlap well, where training continues for 80 seconds after rollout completes.
5.1.2 Embodied RL Training
Figure 14: End-to-end throughput of RLinf and SimpleVLA-RL under different cluster scales.
Experimental setup.   We evaluate on OpenVLA [kim24openvla] and OpenVLA-OFT [kim2025fine], two vision-language-action models that are supervised-finetuned using RL4VLA [liu2025what] and SimpleVLA-RL [simplevla], respectively. We train OpenVLA on ManiSkill [taomaniskill3] and OpenVLA-OFT on LIBERO [liu2023libero], two widely used embodied environments that emulate physical tasks such as pick-and-place. On LIBERO, we compare RLinf with SimpleVLA-RL (commit d001d) [simplevla], which is built on veRL. On ManiSkill, no distributed RL baseline exists, so we compare different execution modes of RLinf. Training speed is reported in steps/sec, computed as the total number of environment steps divided by the iteration time.
ManiSkill environment.   We train OpenVLA on the “PutCarrotOnPlateInScene-v2” task [liu2025what] using 256 parallel ManiSkill environments, each stepping for 80 steps per iteration. Figure 14(a) reports end-to-end throughput across three execution modes in RLinf on 8, 16, and 32 GPUs. RLinf-Hybrid achieves 52.2–69.1% higher throughput than RLinf-Temporal and 60.7–87.2% higher than RLinf-Spatial. For ManiSkill environment, the per-step environment time remains nearly constant as parallelism increases (3(b)) and scaling parallel environments is primarily limited by GPU memory. Thus, dedicating GPUs to environment execution is advantageous. As rollout (i.e., environment and generation) and training all require GPU memory, temporally sharing GPUs between rollout and training is more effective. Thus, RLinf-Hybrid performs much better than the other two. In RLinf-Temporal, environments and generation must coexist in memory, constraining both the number of environments and the batch size, ultimately reducing GPU utilization. RLinf-Spatial splits GPUs across the components, forcing training to retain a sufficient number of GPUs to be runnable, which leaves too few GPUs for rollout and significantly slows it down.
Figure 15: Latency breakdown of ManiSkill and LIBERO.
LIBERO environment.   We train OpenVLA-OFT on the public benchmark task groups provided by LIBERO, using 512 parallel environments, each running 64 steps per iteration. The results in Figure 14(b) show that RLinf-Temporal is 37.8%, 42.6%, 143.4% faster than SimpleVLA-RL on 8, 16, 32 GPUs respectively. As shown in Figure 15, both rollout and training are faster in RLinf-Temporal than in SimpleVLA-RL due to (1) the elimination of redundant environment initialization during rollout, and (2) the use of a single forward pass to compute both action and log probability, with only a modest memory increase. RLinf-Temporal also scales better thanks to its decoupled system design and more efficient implementation. In contrast to ManiSkill, RLinf-Hybrid performs worse than RLinf-Temporal because LIBERO environment is CPU-intensive; allocating environments on a subset of GPUs limits the utilization of CPU cores, making its rollout stage even slower than that of SimpleVLA-RL.
Figure 16: (a) The real and predicted latency in different cases. (b) The search overhead under different cluster scales.
5.2 Effectiveness of Search Policy
We evaluate the search policy along two dimensions: (a) the accuracy of end-to-end throughput estimates derived from profiling data, and (b) the search speed as the number of GPUs increases.
Estimation accuracy.   Figure 16(a) compares real and predicted iteration times across four Qwen2.5–GRPO settings. Cases 1 and 2 use the 1.5B model on 128 GPUs with temperatures 0.6 and 1.0, while Cases 3 and 4 use the 7B model on 64 GPUs with the same temperature settings. Temporal mode achieves estimation errors below 2% because its end-to-end time is well approximated by summing profiled worker times. Spatial mode with pipelining shows slightly higher errors (<5%), mainly due to response-length variability that causes pipeline imbalance. These errors are small enough that they do not change the ranking of feasible execution modes, allowing the search policy to consistently select the most efficient mode for the above end-to-end experiments.
Search speed.   Search complexity is dominated by the number of workflow-graph nodes and available GPUs. RL workflows typically have fewer than ten nodes, allowing our dynamic-programming search to finish quickly. For a graph with three nodes (e.g., Qwen2.5-GRPO or ManiSkill) and GPU counts ranging from 8 to 1024 (Figure 16(b)), search time grows exponentially with GPU count but remains under five seconds, demonstrating the efficiency of our algorithm.
5.3 Model Performance
We comprehensively evaluate the performance of the models produced by the reasoning and the embodied RL training, respectively, to demonstrate the effectiveness of RLinf.
@@ 5.3 Model Performance
Table 1: Evaluation scores of RLinf-math-1.5B/7B and open-source models. GPQA represents GPQA-diamond.
Model AIME24 AIME25 GPQA Avg.
@@
Base model 28.33 24.90 27.45 26.89
DeepScaleR [deepscaler2025] 40.41 30.93 27.54 32.96
AReaL [areal] 44.42 34.27 33.81 37.50
FastCuRL [fastcurl] 43.65 32.49 35.00 37.05
RLinf 48.44 35.63 38.46 40.84
@@
Base model 54.90 40.20 45.48 46.86
AReaL [areal] 61.66 49.38 46.93 52.66
Skywork [skywork-or1-2025] 66.87 52.49 44.43 54.60
Polaris [Polaris2025] 68.55 51.24 43.88 54.56
AceMath [acemath2024] 67.30 55.00 45.57 55.96
RLinf 68.33 52.19 48.18 56.23
Table 2: OpenVLA model success rate results on ManiSkill3.
Model Vision Semantic Position Avg.
RL4VLA [liu2025what] 80.47% 75.00% 81.77% 79.15%
RLinf 82.03% 78.35% 85.42% 81.93%
Reasoning RL Training.   We applied GRPO training, a built-in RL algorithm in RLinf, on two base models, i.e., DeepSeek-R1-Distill-Qwen-1.5B and DeepSeek-R1-Distill-Qwen-7B, to improve their math reasoning ability. As shown in Table 1, the models trained with RLinf achieve the best average performance compared to the baseline systems. RLinf 1.5B model outperforms the baselines across all the three benchmarks (i.e., AIME 24 [aime24], AIME 25 [aime25], GPQA-diamond [rein2024gpqa]), improving by up to 20 points over its base model. RLinf 7B model achieves the highest performance on GPQA-diamond.
Table 3: OpenVLA-OFT success rate after RL on LIBERO.
Model Spatial Object Goal Long Avg.
OpenVLA-OFT(one-traj)  [simplevla] 56.45% 25.60% 45.59% 9.68% 34.33%
RLinf 98.99% 98.99% 98.99% 94.35% 97.83%
Embodied RL Training.   We also evaluate the performance of the models after embodied RL training using RLinf. Table 2 and Table 3 present the evaluation results on the ManiSkill and LIBERO tasks by training OpenVLA and OpenVLA-OFT models respectively, using the PPO algorithm. For ManiSkill, the trained OpenVLA achieves higher success rate than the RL4VLA [liu2025what] baseline system. For LIBERO, we further train the publicly released OpenVLA-OFT model (one-trajectory fine-tuned version) from SimpleVLA-RL [simplevla]. By comparing its task success rates with those of the model RL trained using RLinf, we find that RLinf significantly improves OpenVLA-OFT’s performance on the LIBERO tasks.
@@ 6 Related Works
RL Training Frameworks.   RL training frameworks adopt varying system designs to support large-scale alignment tasks [rlhf, deeprl], typically falling into either task-colocated or task-separated execution modes. For task-colocated systems, DeepSpeed Chat [deepspeedchat] and veRL[verl] put all training phases on shared GPUs for simplified orchestration. In contrast, task-separated systems like NeMo-Aligner [nemo] and OpenRLHF [openrlhf] divide components across devices to improve modularity and scalability. AReal [areal] further introduces asynchronous model updates algorithm upon task-separated systems to increase training throughput. Unlike these frameworks, RLinf provides a more flexible component-to-device placement mechanism, enabling practitioners to explore and deploy more efficient execution configurations tailored to workload characteristics.
Distributed Training Systems for LLM.   Distributed training frameworks for LLMs [fsdp, alpa, colossal-ai] and system-level optimizations for memory [recompute, offload, stalloc] and communication [centauri, syndicate, p3] have significantly advanced large-model scaling. Megatron-LM [megatron] combines tensor, pipeline, and data parallelism to improve scalability, while DeepSpeed [deepspeed] introduces the ZeRO family of techniques [zero, zero-offload, zero-infinity] that shard optimizer states, gradients, and activations to reduce GPU memory usage. Although distributed RL systems face similar multi-device scaling challenges, RL workflows are more dynamic due to interactive data generation and asynchronous updates, making system design and optimization substantially more complex.
Dataflow System.   Traditional dataflow systems [spark, dryad, naiad, mapreduce] excel at large-scale data processing through static task graphs and centralized scheduling, making them efficient for batch and streaming workloads with predictable structures. In contrast, RL training pipelines feature dynamic task graphs and asynchronous components (e.g., data collection, policy updates), requiring more flexible coordination than existing systems. Modern frameworks such as Ray [ray] fill this gap with actor-based execution and decentralized control, making them better suited for orchestrating complex RL workflows.